# Refining Salience-Aware Sparse Fine-Tuning Strategies for Language Models

### v1

@@ 4.1 Main Results
Our experiments results on OPT-350m and BERT-base-uncased can be seen in Table 1. For additional results on RoBERTa-base, OPT-125m and OPT-1.3b, please refer to Tables 8, 9 and 10 in Appendix B. Across all models, we observed that among all the approaches, gradient-based SPEFT has the best average accuracy, higher than LoRA and PiSSA. For instance, in OPT-125m and OPT-350m, gradient-based SPEFT achieves $86.92\%$ and $88.45\%$, that are higher than the best competing SPEFT methods by $0.73\%$ and $0.85\%$ respectively. Particularly on OPT-350m, gradient-based SPEFT has the best performance on MNLI, MRPC, SST-2, and STS-B, On QNLI and QQP, LoRA has the best performance while gradient-based SPEFT has a good performance close to it. This shows that although LoRA shows excellent performance on certain tasks, SPEFT methods, particularly with the gradient salience metric, could further push the limit, achieving better results in accuracy. On BERT-base-uncased, we found that while SPEFT with Fisher-Info salience metric outperforms gradient-based SPEFT on QNLI, QQP and SST-2, it has a large gap in performance in the remaining tasks, making gradient-based SPEFT a more reliable and desirable choice. Similar results are also observed for other OPT variants in Tables 9 and 10 and RoBERTa-base in Table 8 of Appendix B.
@@ 4.2 Larger Scale Models
To evaluate on the text generation task, We fine-tuned Gemma2-2B with our methods We also provide the results of the pretrained model (without fine-tuning) and LoRA as baselines. The results are shown in Table 3. It can be seen that the sparse adapters outperformed the LoRA baseline, with the gradient-based SPEFT method leading the pack with the best performance. Notably, the lead by sparse adapters widens as the task complexity increases, which demands token sequence generation with multi-step reasoning.
Table 3: Comparing the salience metrics on Gemma2-2B with 0.97% trainable parameters. We fine-tuned the model on MetaMathQA and evaluated on 8-shot GSM8K. For reference, we provide pretrained model (without fine-tuning) and the LoRA baseline with the same number of trainable parameters.
@@ 4.3 Exploration of masking strategies
Based on the comparisons with SPEFT in Section 4.1, which showed that gradient-based SPEFT is the best-performing method, we would use it for ablation studies of dynamic vs.missing static masks, and global vs.missing local sparsity. In this section, we delve into the comparisons between between global and local sparsity (Section 3.3) and also static and dynamic masking strategies (Section 3.4) using gradient-based SPEFT, the best-performing salience metric, across OPT-125m, OPT-350m, and BERT-base-uncased. Here, we periodically update the masks every $I=1000$ steps with 1024 training examples to estimate the salience metrics. The results are shown in Table 4.
@@ 5 Discussion
Numerous hardware vendors have introduced specialized hardware features with instruction set extensions tailored for sparse matrix multiplication. Especially in recently announced hardware devices. Mainstream devices like NVIDIA’s A100 [6], H100 [5], and H200, as well as offerings from other major vendors or emerging competitors such as AMD’s MI300 [1] and Cerebras’ WSE2 [27], are embracing this trend. As hardware support for sparse computation advances, the utility of sparsity-based PEFT, or generally sparse training, is poised to improve substantially. This development will enable both current and future strategies to attain performance levels closer to their full potential, as these calculations won’t require emulation via dense computations, allowing for closer realization of theoretical speedups and savings on FLOPs.
@@
A fundamental element of this study involves reevaluating certain design choices in SPEFT, leading to the discovery that straightforward designs, such as first-order salience proxies, emerge as the most effective methods. Intriguingly, selecting the most salient weights in a neural network has being a long-standing challenge, one that dates back to early weight pruning research by LeCun et al.missing in 1989 [19]. It’s notable that the optimal saliency metric seems to differ – or arguably should differ – among different task setups, such as post-training weight pruning [19], pruning at initialization [21, 9], and zero-cost NAS proxies [28]. The suggested practice then should be to systematically review a range of known and established proxies to set a solid baseline before designing a complex salience metric.
@@ 7 Limitations
8 Computational Resources
@@
Appendix A Hyperparameters
The hyperparameters we used for all models are shown in Tables 7, 6 and 5. Notably, for all models, the density ${\rho}$ was set to make sure the number of trainable parameters across all methods was the same as the LoRA baseline.
Table 5: The hyperparameters we used for Gemma2-2B on MetaMathQA.
Model Method Hyperparameters MetaMathQA Optimizer AdamW Shared Warmup Ratio 0.03 LR Schedule Linear Gemma2-2B (LoRA) Batch Size 16 # Epochs 1 Learning Rate 2E-5 LoRA $r$ 64 LoRA $\alpha$ 16 Max Seq. Len. 1024 Gemma2-2B (Sparse) Batch Size 16 # Epochs 1 Learning Rate 2E-5 Sparse Top $k$ 0.18% Max Seq. Len. 1024
Method Dataset MRPC, STS-B QNLI, SST-2, MNLI, QQP Shared Optimizer AdamW AdamW Warmup Ratio 0 0 LR Schedule Linear Linear Batch Size 16 64 # Epochs 30 30 Learning Rate 4E-4 5E-5 Max Seq. Len. 512 196 LoRA LoRA $r$ 8 8 LoRA $\alpha$ 16 8 OPT-125m Sparse ${\rho}$ 0.35% 0.35% OPT-350m Sparse ${\rho}$ 0.35% 0.35% BERT-base Sparse ${\rho}$ 0.27% 0.27% RoBERTa-base Sparse ${\rho}$ 0.24% 0.24% Model (Method) Hyperparameters All datasets Shared Optimizer AdamW Warmup Ratio 0 LR Schedule Linear Learning Rate 5E-5 # Epochs 30 Batch Size 16 OPT-1.3b (LoRA) LoRA $r$ 8 LoRA $\alpha$ 8 OPT-1.3b (Sparse) Sparse ${\rho}$ 0.18%
Appendix B Additional Experimental Results
Tables 8, 9 and 10 provide additional respective results on GLUE tasks for the OPT-125m and OPT-1.3b variants, and BERT-base-uncased.

### v2

@@ 4.1 Main Results
Our experiments results on OPT-350m and BERT-base-uncased can be seen in Table 1. For additional results on RoBERTa-base, OPT-125m and OPT-1.3b, please refer to Tables 9, 10 and 11 in Appendix B. Across all models, we observed that among all the approaches, gradient-based SPEFT has the best average accuracy, higher than LoRA and PiSSA. For instance, in OPT-125m and OPT-350m, gradient-based SPEFT achieves $86.92\%$ and $88.45\%$, that are higher than the best competing SPEFT methods by $0.73\%$ and $0.85\%$ respectively. Particularly on OPT-350m, gradient-based SPEFT has the best performance on MNLI, MRPC, SST-2, and STS-B, On QNLI and QQP, LoRA has the best performance while gradient-based SPEFT has a good performance close to it. This shows that although LoRA shows excellent performance on certain tasks, SPEFT methods, particularly with the gradient salience metric, could further push the limit, achieving better results in accuracy. On BERT-base-uncased, we found that while SPEFT with Fisher-Info salience metric outperforms gradient-based SPEFT on QNLI, QQP and SST-2, it has a large gap in performance in the remaining tasks, making gradient-based SPEFT a more reliable and desirable choice. Similar results are also observed for other OPT variants in Tables 10 and 11 and RoBERTa-base in Table 9 of Appendix B.
@@ 4.2 Larger Scale Models
To evaluate on the text generation task, We fine-tuned Gemma2-2b with our methods on MetaMathQA and evaluated on 5-shot GSM8K. We also provide the results of the pretrained model (without fine-tuning) and LoRA as baselines. The results are shown in Table 3. It can be seen that the sparse adapters outperformed the LoRA baseline, with the gradient-based SPEFT method leading the pack with the best performance. Furthermore, for code generation tasks, we fine-tuned Llama3-8b with our methods and evaluated on HumanEval and MBPP benchmarks. The results are shown in Table 12 of Appendix B. Notably, the lead by sparse adapters widens as the task complexity increases, which demands token sequence generation with multi-step reasoning.
@@
Table 3: Comparing the salience metrics on Gemma2-2b with 0.97% trainable parameters. We fine-tuned the model on MetaMathQA and evaluated on 5-shot GSM8K. For reference, we provide pretrained model (without fine-tuning) and the LoRA baseline with the same number of trainable parameters.
@@ 4.3 Exploration of masking strategies
Based on the comparisons with SPEFT in Section 4.1, which showed that gradient-based SPEFT is the best-performing method, we would use it for ablation studies of dynamic vs.missing static masks, and global vs.missing local sparsity. In this section, we delve into the comparisons between global and local sparsity (Section 3.3) and also static and dynamic masking strategies (Section 3.4) using gradient-based SPEFT, the best-performing salience metric, across OPT-125m, OPT-350m, and BERT-base-uncased. Here, we periodically update the masks every $I=1000$ steps with 1024 training examples to estimate the salience metrics. The results are shown in Table 4.
@@ 5 Discussion
Numerous hardware vendors have introduced specialized hardware features with instruction set extensions tailored for sparse matrix multiplication. Especially in recently announced hardware devices. Mainstream devices like NVIDIA’s A100 Choquette et al. (2021), H100 Choquette (2023), and H200, as well as offerings from other major vendors or emerging competitors such as AMD’s MI300 AMD and Cerebras’ WSE2 Selig (2022), are embracing this trend. As hardware support for sparse computation advances, the utility of sparsity-based PEFT, or generally sparse training, is poised to improve substantially. This development will enable both current and future strategies to attain performance levels closer to their full potential, as these calculations won’t require emulation via dense computations, allowing for closer realization of theoretical speedups and savings on FLOPs.
@@
A fundamental element of this study involves reevaluating certain design choices in SPEFT, leading to the discovery that straightforward designs, such as first-order salience proxies, emerge as the most effective methods. Intriguingly, selecting the most salient weights in a neural network has being a long-standing challenge, one that dates back to early weight pruning research by LeCun et al.missing in 1989 LeCun et al. (1989). It’s notable that the optimal saliency metric seems to differ – or arguably should differ – among different task setups, such as post-training weight pruning LeCun et al. (1989), pruning at initialization Lee et al. (2019b); de Jorge et al. (2021), and zero-cost NAS proxies Siems et al. (2020). The suggested practice then should be to systematically review a range of known and established proxies to set a solid baseline before designing a complex salience metric.
@@ 6 Conclusion
7 Acknowledgments
This work is supported in part by the National Key R&D Program of China (2023YFC3321600), National Natural Science Foundation of China (62376263, 62372443 and 62271496), Guangdong Basic and Applied Basic Research Foundation (2023B1515130002), Natural Science Foundation of Guangdong (2024A1515030209 and 2024A1515011970), Shenzhen Science and Technology Innovation Commission (JCYJ20230807140507015 and JCYJ20220531100804009).
@@ 8 Limitations
Appendix A Hyperparameters
The hyperparameters we used for all models are shown in Tables 5, 6, 8 and 8. Notably, for all models, the density ${\rho}$ was set to make sure the number of trainable parameters across all methods was the same as the LoRA baseline.
Method Dataset MRPC STS-B QNLI SST-2 MNLI QQP Shared Optimizer AdamW AdamW Warmup Ratio 0 0 LR Schedule Linear Linear Batch Size 16 64 # Epochs 30 30 Learning Rate 4E-4 5E-5 Max Seq. Len. 512 196 LoRA LoRA $r$ 8 8 LoRA $\alpha$ 16 8 OPT-125m Sparse ${\rho}$ 0.35% 0.35% OPT-350m Sparse ${\rho}$ 0.35% 0.35% BERT-base Sparse ${\rho}$ 0.27% 0.27% RoBERTa-base Sparse ${\rho}$ 0.24% 0.24% Model (Method) Hyperparameters All datasets Shared Optimizer AdamW Warmup Ratio 0 LR Schedule Linear Learning Rate 5E-5 # Epochs 30 Batch Size 16 OPT-1.3b LoRA $r$ 8 (LoRA) LoRA $\alpha$ 8 OPT-1.3b (Sparse) Sparse ${\rho}$ 0.18%
Model (Method) Hyperparameters MetaMathQA Optimizer AdamW Shared Warmup Ratio 0.03 LR Schedule Linear Gemma2-2b (LoRA) Batch Size 16 # Epochs 1 Learning Rate 2E-5 LoRA $r$ 64 LoRA $\alpha$ 16 Max Seq. Len. 1024 Gemma2-2b (Sparse) Batch Size 16 # Epochs 1 Learning Rate 2E-5 Sparse Top $k$ 0.18% Max Seq. Len. 1024 Model (Method) Hyperparameters CodeFeedback Optimizer AdamW Shared Warmup Ratio 0.03 LR Schedule Linear Llama3-8b (LoRA) Batch Size 16 # Epochs 1 Learning Rate 2E-5 LoRA $r$ 64 LoRA $\alpha$ 16 Max Seq. Len. 512 Llama3-8b (Sparse) Batch Size 16 # Epochs 1 Learning Rate 2E-5 Sparse Top $k$ 0.67% Max Seq. Len. 512
Table 7: The hyperparameters we used for Gemma2-2b on MetaMathQA. The percentage of trainable parameters (${\rho}$) for the sparse models are chosen to be the same as the LoRA models.
Table 8: The hyperparameters we used for Llama3-8b on CodeFeedback. The percentage of trainable parameters (${\rho}$) for the sparse models are chosen to be the same as the LoRA models.
Method HumanEval MBPP Avg. LoRA 40.85 48.8 44.83 PiSSA 38.41 48.0 43.21 Gradient 48.78 50.0 49.39 SynFlow 40.85 49.0 44.93 Magnitude 39.02 49.2 44.11 SNIP 46.95 51.0 48.98 FORCE 46.95 50.4 34.91 Taylor-FO 49.39 49.4 49.40 GRaSP 46.34 49.8 48.07 Fisher-Info 46.34 47.2 46.77
Table 12: Comparing the salience metrics on Llama3-8b with 0.67% trainable parameters. We fine-tuned the model on CodeFeedback and evaluated on HumanEval and MBPP. For reference, we provide LoRA and PiSSA as baselines with the same number of trainable parameters.
Appendix B Additional Experimental Results
Tables 9, 10 and 11 provide additional respective results on GLUE tasks for the OPT-125m and OPT-1.3b variants, and BERT-base-uncased. Table 12 shows the results on HumanEval and MBPP benchmarks for Llama3-8b model.
Appendix C Additional Ablation Studies
We fine-tuned the Gemma2-2b model on the MetaMathQA dataset and evaluated it on the GSM8K_cot task (5-shot) using flexible extract and strict match metrics. In order to explore the efficiency-performance trade-off, we varied for LoRA $r$ from 4 to 128 and compare it against SPEFT methods with the same trainable parameters for each config. The LoRA $\alpha$ was always kept the same as $r$.
With the same numbers of training parameters, LoRA and SPEFT would use almost identical FLOPs per step, as the added overheads of both are of the same magnitude and much smaller (<0.5% in all of our main experiments) than the base model. There was no noticeable difference between LoRA and SPEFT in terms of computational and memory footprint for all runs.
As is shown in Figure 2, the performance of SPEFT methods improve with increasing trainable parameters while LoRA results are mostly constant with increased parameter budget. Overall, the gradient-based SPEFT outperformed LoRA using fewer trainable parameters, but also widens the gap further as the budget increases.
Appendix D Computational Resources
@@
(a) Flexible Extract.
(b) Strict Match.
Figure 2: Varing the number of trainable parameters on Gemma2-2b and GSM8K_cot (5-shot) with LoRA, PiSSA and SPEFT methods. The x-axis represents the percentage of trainable parameters, while the y-axis denotes accuracy.